OpenStudy (anonymous):

A Poisson process can be used to represent the occurrence of structural loads over time. Suppose the mean time between occurrences of loads is 0.60 year. (c) How long must a time period be so that the probability of no loads occurring during that period is at most 0.13?

OpenStudy (mathmate):

Since it is a Poisson process, we can safely use the Poisson distribution to model the problem. Given \(\lambda=0.6 \)= mean time between occurrences, and the Poisson distribution (pdf) is given by: P(X=k)=\(\dfrac{e^{-\lambda}\lambda^k}{k!}\) We set k=0 (no occurrence) and P(X=0)=0.13 to solve for lambda. 0.13=\(\dfrac{e^{-\lambda}\lambda^0}{0!}=e^{-\lambda}\) Solving gives \(\lambda=2.04\) which means the time period would have an average of 2.04 occurrences, or time = 2.04*0.6 = 1.224 years.

OpenStudy (anonymous):

thanks a buch!!

OpenStudy (mathmate):

You're welcome! :)